OpenStudy (anonymous):

Could someone explain how to do this problem? Capacitors of 10mF and 20mF are connected in series, and this pair is then connected in parallel with a 30mF CAPACITOR. What is the equivalent capacitance of this arrangement?

OpenStudy (anonymous):

just go to hyperphysics, capacitor button. series parallel combination. the formula is there.

OpenStudy (anonymous):

Is it (1/10)+(1/20)+(30)?

OpenStudy (anonymous):

For capacitance in series the formula is \[\frac{ 1 }{ Ceq } = \frac{ 1 }{ C _{1} } + \frac{ 1 }{ C _{2} } + . . . \frac{ 1 }{ C _{n} }\] and for parallel it's simply \[Ceq = C _{1} + C _{2} +. . . C _{n}\] so your equation would be \[C _{12}= \frac{ 1 }{ 10 } + \frac{ 1 }{ 20 } \] then \[C _{12} + 30 mF\] since in parallel

OpenStudy (whpalmer4):

capacitors in series combine with the same formula as resistors in parallel. capacitors in parallel combine with the same formula as resistors in series.